OpenStudy (anonymous):

What is the standard form equation of the line passing through (–2, 8) with a slope of 2? I am having trouble figuring out how to do this problem... I think I was absent this day Below is the choices A. 2x – y = 12 B. 2x – y = –12 C. 2x + y = 12 D.2x + y = –12

OpenStudy (anonymous):

The slope-intercept formula is all you need: \[y-y _{1} = m(x-x _{1})\] The x1 and y1 are the two values (-2, 8) you were given, and the m is the slope you were also given. Plug and chug.

OpenStudy (anonymous):

so y-8=2(x+2)?

OpenStudy (anonymous):

Exactly, so just distribute the 2 through the parentheses then rearrange the equation so you have the x's and y's on the left side and the 12 on the right side like in your choices of solutions.

OpenStudy (zepp):

I have a tutorial on this thing http://openstudy.com/users/zepp#/updates/4fa3474fe4b029e9dc34125e You'd have to scroll down for those cases.

OpenStudy (zepp):

Situation 3 :)

OpenStudy (anonymous):

so... y-8=2x+4 -2x+y=12 so..... -1(-2x+y=12) 2x-y=-12 sooo B

OpenStudy (anonymous):

Oh it's called the "point-slope form" as I see from your tutorial zepp, not the "slope-intercept" form as I was calling it. My bad :) But yes lyssa, that is correct! :)

OpenStudy (anonymous):

YAY:D Thanks I got some more question that I will be asking :D

OpenStudy (zepp):

Yeah, slope-intercept form is \(y=mx+b\)

OpenStudy (anonymous):

Can I ask here?

OpenStudy (anonymous):

Or to cut down on your steps a little, you don't have to multiply by (-1). You can just do: y-8=2x+4 -8-4 = 2x-y -12=2x-y To get the same answer. If your other questions are related to this one then I assume you would be fine in posting them here but I am a newbie to this website so I do not know.

OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)?

OpenStudy (zepp):

http://openstudy.com/users/zepp#/updates/4fa3474fe4b029e9dc34125e Situation 4 :)

OpenStudy (anonymous):

Thanks Zepp:D

OpenStudy (zepp):

You are welcome, you'll get a point-intercept form when you simplify everything, so all you need to do is to transform it back into the point-slope form.

OpenStudy (anonymous):

It wont go to the site so can you help me here

OpenStudy (zepp):

http://puu.sh/F7Lc There you go

OpenStudy (anonymous):

(–2, 0) and (2, 8) m=8-0/2+2=2 so slope is 2

OpenStudy (zepp):

You are on the right track :)

OpenStudy (anonymous):

y-y1=m(x-x1) y-0=2(x+2) y-0=2x+4

OpenStudy (zepp):

You are so goooood! :D

OpenStudy (anonymous):

2x-y=-4

OpenStudy (anonymous):

these are my choices y = 2(x + 2) y = 2(x – 2) y = –2(x – 2) y = –2(x + 2)

OpenStudy (anonymous):

sooo. A